Let U be the subspace of R 2 spanned by (1, 2). Find the orthogonal complement U ⊥ of U. Then find a ∈ U and b ∈ U ⊥ such that (0, 3) = a + b.
Answer #1

P (12) Lsay 7 CER Solution : Orthogonal complement utef a subspace. U is given by utzfeul Teus 20 vueug @ Given that .. Visa:: (0,3) = atb, so cand. Hence la =(0,3)+c (1,4/2) EU $16= -2 (1-7) Eut